# Deep Unlearn: Benchmarking Machine Unlearning

### v1

@@ 3.2 State-of-the-art MU methods
Prune-Reinitialize-Match-Quantize (PRMQ) [4] first prunes the model via L1 pruning, reinitializes parts of the model, optimises it using a combination of cross-entropy and a mean-squared-error on the entropy between the outputs of $f_{O}$ and $f_{U}$ on $\mathcal{D}_{R}$ and finally converts $f_{U}$’s weights to half-precision floats.
@@
Knowledge-Distillation-Entropy (KDE) [6] uses a teacher-student setup. Both student and teacher start as copies of the original model, then the student’s first and last layers are re-initialised. The student $f_{U}$ minimizes its Kullback-Leibler Divergence (KLD) with the $f_{O}$ over $\mathcal{D}_{V}$, then minimizes a combination of losses: a soft cross-entropy loss between $f_{U}$ and $f_{O}$, a cross-entropy loss on outputs of $\mathcal{D}_{R}$ from $f_{U}$, and the KLD between $f_{U}$ and $f_{O}$ on $\mathcal{D}_{R}$.
Repeated-Noise-Injection (RNI) [7] first reinitialises the final layer of the model, then repeatedly injects noise in different layer of the model while training on the $\mathcal{D}_{R}$.
@@
Exact Unlearning - K (EU-K) [24] freezes the first layers then restores the weights of the last $k$ layers to their initialization state. We randomly reinitialize the weights instead, so that the method no longer requires knowledge about the training process of $f_{O}$.
SCRUB [37] leverages a student-teacher setup where the model is optimised for three objectives: matching the teacher’s output distribution on $\mathcal{D}_{R}$, correctly predicting the $\mathcal{D}_{R}$ set and ensuring the output distributions of the teacher and student diverge on the $\mathcal{D}_{F}$
@@
Bad Teacher (BT) [16] uses a teacher-student approach with two teachers: the original model, and a randomly initialized model - the bad teacher-, the student starts as copy of $f_{U}$ then learns to mimic the $f_{O}$ on $\mathcal{D}_{R}$ and the bad teacher on the $\mathcal{D}_{F}$.
@@ 4 Experimental Evaluation
A further observation is that prior research tends to compare MU methods with default hyperparameters, potentially leading to a less competitive performance of the method. To ensure that we get the best performance out of each method, we perform three hyperparameter sweeps to find the best set of hyperparameters for each method. To ensure a fair comparison, we use same number of searches for each method. Each hyper-parameter sweep uses 100 trials to minimize four loss functions: Retain Loss ($\mathcal{L}_{\text{Retain}}$), Forget Loss ($\mathcal{L}_{\text{Forget}}$), Val Loss ($\mathcal{L}_{Val}$), and Val MIA ($\mathcal{L}_{\text{Val-MIA}}$) given by
$\mathcal{L}_{Retain}=\alpha\times|\text{RA}(f_{U})-\text{RA}(f_{R})|,\;\mathcal{L}_{Forget}=\beta\times|\text{FA}(f_{U})-\text{FA}(f_{R})|,$ (5)
$\mathcal{L}_{Val}=\gamma\times|\text{VA}(f_{U})-\text{VA}(f_{R})|,\;\mathcal{L}_{\text{Val-MIA}}=\eta\times\text{Disc}(\mathcal{D}_{V},f_{U}),$ (6)
@@
where the $\mathcal{L}_{\text{Retain}}$ captures the divergence in accuracy between the retrained and unlearned model over the $\mathcal{D}_{R}$, $\mathcal{L}_{\text{Forget}}$ and $\mathcal{L}_{Val}$ capture the divergence over $\mathcal{D}_{F},\mathcal{D}_{V}$ respectively and $\mathcal{L}_{\text{Val-MIA}}$ captures whether the loss distributions over $\mathcal{D}_{F}$ and $\mathcal{D}_{V}$ are distinguishable from one another via the discernibility score defined in Section 2. We set $\alpha=\beta=\gamma=\frac{1}{3}$ and $\eta=1$ as we found these values to balance the importance of importance retention and the resilience to Membership Inference Attacks. Per unlearn method, we use the hyperparameter configuration that minimises the four loss terms when evaluating the method. Thus, for each unlearning method, we first unlearn $300$ models to do the hyper-parameter sweep, then unlearn $10$ models with the best set of hyper-parameters, leading to $5,580$ per dataset for a given architecture, leading to a total of $50,220$ for the 9 dataset / model combinations.
@@
Ranking. A challenge in the comparison of MU method performance comes from the potential proximity of the evaluation metrics. As a simple example, suppose we have four methods $U_{1},...,U_{4}$ with accuracies: 98%, 99%, 50%, 1%, respectively; if we simply rank the methods, the rank itself would not be representative of the fact that e.g. $U_{1}$ and $U_{2}$ are much above $U_{3}$ and $U_{4}$. In order to enable distinctions based on proximities, we use Agglomerative Clustering and define cut-off points such that we obtain three clusters: (1) Best performers (G1), (2) Average performers (G2), and (3) Worst performers (G3). If a method does not produce 10 usable models, one per original model, it is assigned to a Failed group (F). For each method, we count the number of times it appears in each of the three groups (with nine being the maximum). To obtain a final ranking of the methods, we first rank the methods using the number of times it appears in the Best Performers group (G1); if ties occur, we use the Average Performers (G2) group to break them. If ties persist, the Worst Performers (G3) group serves as the final tie-breaker. This method ensures a clear and fair ranking by considering each performance group in order of importance.
@@ 5 Main Results
Table 1 presents the main results of our evaluations on MU methods based on Retention Deviation and Indiscernibility. The results for the run-time efficiency are shown in Table 2.
Table 1: Ranking by performance on Retention Deviation and Indiscernibility across datasets and architectures. We count the number of times each method appears in the Best Performers group (G1), Average performance group (G2) and Worst performers group (G3) (see §4). The final rank is computed based on the number of times the method appears in G1—with occurrences in G2 and G3 used to break ties if needed. If a method does not produce any usable models, it is assigned to a Failed group (F). Three methods appear in the top 3 for both performance measures: MSG (1st and 1st), CT (3rd and 1st) and KDE (3rd and 2nd).
Retention Deviation Indiscernibility
Rank Method G1 G2 G3 F Rank Method G1 G2 G3 F
1 FT 8 1 0 0 1 CT 9 0 0 0
1 MSG 8 1 0 0 1 MSG 9 0 0 0
2 PRMQ 7 2 0 0 2 CFW 7 2 0 0
3 CT 7 1 1 0 2 RNI 7 2 0 0
3 KDE 7 1 1 0 2 KDE 7 2 0 0
3 CFW 7 1 1 0 3 FT 6 3 0 0
4 FCS 6 3 0 0 3 PRMQ 6 3 0 0
4 SalUN 6 3 0 0 3 SalUN 6 3 0 0
5 NG+ 5 4 0 0 4 SRL 6 2 1 0
5 SRL 5 4 0 0 5 NG+ 5 4 0 0
6 SCRUB 4 3 1 1 5 FCS 5 4 0 0
7 BT 2 7 0 0 6 SCRUB 5 3 0 1
7 RNI 2 7 0 0 7 BT 4 5 0 0
8 CF-k 2 3 2 2 8 CF-k 1 2 4 2
9 IU 1 0 2 6 9 EU-k 1 2 2 4
10 EU-k 0 5 0 4 10 GA 0 4 4 1
11 GA 0 1 7 1 11 IU 0 0 3 6
12 FF 0 0 0 9 12 FF 0 0 0 9
On the reliability of baselines.
@@ 12 FF 0 0 0 9 12 FF 0 0 0 9
The commonly used baseline FT trains the original model only on the retain set for several epochs to enable the model to forget information about the forget set. In our evaluation, FT performs best based on the Retention Deviation, and is ranked third based on Indiscernibility. The latter observation may come from the fact that FT does not explicitly unlearn the forget set or perturb the model parameters. Based on these results we conclude it is a reasonable baseline to evaluate against. We however remark that since the mechanism underlying FT (training on the Retain set to maintain performance) is common to many other methods, these methods may inherit its susceptibility to MIA. Another common baseline, GA which performs gradient ascent on the forget set, performs poorly across both metrics. Its more recent variation NG+, which uses an additional retain set correction, ranks fifth for both metrics, making it a more suitable baseline.
@@
Table 2: Run Time Efficiency on ResNet for the top performing methods. CT is the fastest on average, MSG runs up to 5x faster than naive retraining.
@@
On the reliability of newly proposed unlearning methods.
@@
MSG obtains a first rank in both Performance Retention Deviation and Indiscernibility. Its unique approach identifies the parameters in the Convolutional layers that most contribute to the information to be forgotten. This strategy, differing from FT, allows MSG to retain performance from the Retain set while modifying the weights that are more relevant to the Forget set. PRMQ ranks second in performance retention, making it one of the top performers. However, it suffers from the same lower performance in Indiscernibility as FT. PRMQ however does not leverage the Forget set; instead, it performs a form of knowledge distillation by attempting to reproduce the results of the original models on the Retain set. Additionally, during the pruning phase, it reinitializes the weights for the MLP and Convolutional layers. CT ranks third in Performance Retention and first in Indiscernibility. It is interesting to note that CT and MSG are both consistently among the top performers in Indiscernibility, where FT performs poorly.
On the robustness across architectures.
@@
Critical for a good MU method is its ability to generalise across various DNN architectures. We conducted experiments with both ResNet18 and TinyViT. Methods such as CT, despite being tailored to Convolutional Neural Networks (CNN) models, still perform competitively when applied to Vision Transformers. Methods such as CT and MSG, while proposed for CNN layers, work well on Vision Transformer as one can leverage the 2D Convolutions used in Positional Encoders. We provide additional details on the ranking based on architectures in Appendix F.
On the speed of the unlearn methods.
@@
MSG, the best MU candidate runs  7.6x faster than retraining from scratch on UTKFace and  3.3x on FashionMNIST. On average MSG runs  5.3x faster than retraining. The fastest method is CT which achieves a speedup of  17.5x compared to Retraining. This speedup stems from its simple approach: transpose the weights of the convolutional layers.
On the performance when evaluated against a stronger MIA.
From the results above, we note that there exist MU methods that perform fast and reliably across datasets, architectures, and initializations. Specifically, CT, FT, MSG seem strong candidates for reliable MU methods. However, as has also been highlighted in prior work [46], MIA has been debated as a strong metric, as its ability to assess MU is hampered by its own ability to infer data membership. In line with this, recent works have introduced more powerful variations on MIA [51] with [37] proposing the stronger U-LiRA attack for MU. For the best performers in Table 1, we apply the attack setup from [31] where we generate a total of 640 models (with varying train, retain and forget sets) and for each unlearn method perform a hyperparameter sweep to find the best configuration (for details see Appendix D). We determine for each data point its U-LiRA Inference Accuracy and report for each method its average and standard deviation (see Figure 1). From this, we conclude that MSG as well as CT resist U-LiRA attacks. Both MSG and CT thus not only rank first in terms of Performance Retention and Indiscernibility based on U-MIA, but also are robust against a stronger variation of MIA.

### v2

@@ 7 Repeated-Noise-Injection (RNI)
Prune-Reinitialize-Match-Quantize (PRMQ) [31] first prunes the model via L1 pruning, reinitializes parts of the model, optimizes it using a combination of cross-entropy and a mean-squared-error on the entropy between the outputs of $f_{O}$ and $f_{U}$ on $\mathcal{D}_{R}$ and finally converts $f_{U}$’s weights to half-precision floats.
@@
Knowledge-Distillation-Entropy (KDE) [53] uses a teacher-student setup. Both student and teacher start as copies of the original model, and then the student’s first and last layers are reinitialized. The student $f_{U}$ minimizes its Kullback-Leibler Divergence (KLD) with the $f_{O}$ over $\mathcal{D}_{V}$, then minimizes a combination of losses: a soft cross-entropy loss between $f_{U}$ and $f_{O}$, a cross-entropy loss on outputs of $\mathcal{D}_{R}$ from $f_{U}$, and the KLD between $f_{U}$ and $f_{O}$ on $\mathcal{D}_{R}$.
Repeated-Noise-Injection (RNI) [54] first reinitializes the final layer of the model, then repeatedly injects noise in different layers of the model while training on the $\mathcal{D}_{R}$.
@@
Exact Unlearning - K (EU-K) [13] freezes the first layers and then restores the weights of the last $k$ layers to their initialization state. We randomly reinitialize the weights instead so that the method no longer requires knowledge about the training process of $f_{O}$.
SCalable Remembering and Unlearning unBound (SCRUB) [14] leverages a student-teacher setup where the model is optimized for three objectives: matching the teacher’s output distribution on $\mathcal{D}_{R}$, correctly predicting the $\mathcal{D}_{R}$ set and ensuring the output distributions of the teacher and student diverge on the $\mathcal{D}_{F}$
@@
Bad Teacher (BT) [15] uses a teacher-student approach with two teachers: the original model and a randomly initialized model - the bad teacher-, the student starts as a copy of $f_{U}$ then learns to mimic the $f_{O}$ on $\mathcal{D}_{R}$ and the bad teacher on the $\mathcal{D}_{F}$.
@@ 5 Experimental Evaluation
5.1 Data splits
We use a fixed seed $s=123$ to generate the different data splits. If a Training and Test split exists for every dataset, we keep it and refer to the original training split as the development split. Otherwise, we divide the dataset following a $80\%/20\%$ (Development / Test) split. The Development set is then further divided. We use the Development split for training the model, defining the forget and retain set, while we use the Test set to evaluate the final performance of each method. We split the Development into Training $\mathcal{D}$ and Validation $\mathcal{D}_{V}$ splits following a $80\%/20\%$ split. The Training set is used to train the Original model, while the validation set is used for the hyper-parameters searches. We subdivide the Training set into the Retain $\mathcal{D}_{R}$ and Forget $\mathcal{D}_{F}$ sets following a $90\%/10\%$ split. As such, we consider a $10\%$ forgetting budget, which is common in the literature, and draw these $10\%$ from the Training set following a Uniform distribution.
5.2 Datasets
TABLE II: Summary of the datasets used in the benchmark, with split sizes, number of classes per dataset, task and color information.
Dataset Retain size Forget size Validation size Test size Number of classes Task
MNIST 45,900 5,100 9,000 10,000 10 Handwritten digit classification (Grayscale)
FashionMNIST 45,900 5,100 9,000 10,000 10 Fashion item classification (Grayscale)
CIFAR-10 38,250 4,250 7,500 10,000 10 Object classification (RGB)
CIFAR-100 38,250 4,250 7,500 10,000 100 Fine-grained object classification (RGB)
UTKFace 14,508 1,613 2,846 4,741 5 Age classification (RGB)
5.2.1 CIFAR-10
is a widely used dataset in computer vision and machine learning. It comprises 60,000 32x32 color images in 10 different classes, with 6,000 images per class. The dataset is divided into 50,000 training images and 10,000 testing images. CIFAR-10 represents a diverse range of everyday objects, such as airplanes, automobiles, birds, and cats, making it a challenging task for image classification. The simplicity of the images combined with the variety of categories makes CIFAR-10 a suitable dataset to test the efficacy of machine unlearning algorithms in effectively unlearning information without compromising the model’s performance on the remaining data.
5.2.2 CIFAR-100
is a more complex extension of CIFAR-10, containing 100 classes with 600 images per class, split into 500 training images and 100 testing images per class. Each class is labeled with a ”fine” label and grouped into 20 ”coarse” labels, adding another layer of classification difficulty. The increased number of classes and finer granularity make CIFAR-100 an intriguing dataset for machine unlearning benchmarks. It poses a more significant challenge for models to forget specific classes or groups while retaining knowledge of others, thus testing the unlearning algorithms’ precision and effectiveness in handling more granular and complex datasets.
5.2.3 MNIST
is a well-known benchmark in handwritten digit recognition. It comprises 70,000 grayscale images of handwritten digits (0-9), 60,000 used for training, and 10,000 for testing. Each image is 28x28 pixels in size. We consider MNIST due to its simplicity and extensive research and development history. The simplicity of MNIST allows researchers to focus on the fundamental aspects of unlearning techniques without the additional complexity introduced by color or high resolution, providing a clear assessment of the effectiveness of unlearning algorithms in a controlled setting.
5.2.4 Fashion MNIST
is a more challenging replacement for MNIST. It contains 70,000 grayscale images of fashion items in 10 categories: shirts, trousers, and sneakers. Like MNIST, each image is 28x28 pixels, but the increased complexity and variability of clothing items make it a more challenging classification task. Fashion MNIST provides a more realistic and intricate dataset than MNIST, testing the unlearning algorithms’ ability to handle real-world-like variability and ensuring that they can effectively remove learned information while maintaining performance on a moderately complex dataset.
5.2.5 UTKFace
For each dataset, the Original and Retrained models are trained using the same hyper-parameters (provided in Table VII)
5.3 Neural Network Architectures
5.3.1 ResNet: ResNet18
5.3.2 Vision Transformer (ViT): TinyViT
Introduced by Dosovitskiy et al. [61], adapts the transformer architecture to image classification by treating images as sequences of patches. We consider TinyViT from Wu et al. [10], as it is a compact version of ViT designed to be parameter-efficient while maintaining high performance.
@@ 5.4 Hyper-parameters search
A further observation is that prior research tends to compare MU methods with default hyperparameters, potentially leading to a less competitive performance of the method. To ensure that each method performs optimally, we perform three hyperparameter sweeps to find the best set of hyperparameters for each method. We use the same number of searches for each method to ensure a fair comparison. Each hyper-parameter sweep uses 100 trials to minimize four loss functions: Retain Loss ($\mathcal{L}_{\text{Retain}}$ equation 9), Forget Loss ($\mathcal{L}_{\text{Forget}}$ equation 10), Val Loss ($\mathcal{L}_{Val}$ equation 11), and Val MIA ($\mathcal{L}_{\text{Val-MIA}}$ equation 12) given by
$\mathcal{L}_{Retain}=\alpha\times|\text{RA}(f_{U})-\text{RA}(f_{R})|$ (9)
$\mathcal{L}_{Forget}=\beta\times|\text{FA}(f_{U})-\text{FA}(f_{R})|$ (10)
$\mathcal{L}_{Val}=\gamma\times|\text{VA}(f_{U})-\text{VA}(f_{R})|$ (11)
$\mathcal{L}_{\text{Val-MIA}}=\eta\times\text{Disc}(\mathcal{D}_{V},f_{U})$ (12)
@@
Where the $\mathcal{L}_{\text{Retain}}$ captures the divergence in accuracy between the retrained and unlearned model over the $\mathcal{D}_{R}$, $\mathcal{L}_{\text{Forget}}$ and $\mathcal{L}_{Val}$ capture the divergence over $\mathcal{D}_{F},\mathcal{D}_{V}$ respectively and $\mathcal{L}_{\text{Val-MIA}}$ captures whether the loss distributions over $\mathcal{D}_{F}$ and $\mathcal{D}_{V}$ are distinguishable from one another via the discernibility score defined in Section 3.2.1. We set $\alpha=\beta=\gamma=\frac{1}{3}$ and $\eta=1$ as we found these values to balance the importance of importance retention and resilience to Membership Inference Attacks. Per unlearning method, we use the hyperparameter configuration that minimizes the four loss terms when evaluating the method. Thus, for each unlearning method, we first unlearn $300$ models to do the hyper-parameter sweep and then unlearn $10$ models with the best set of hyper-parameters, which amounts to $5,580$ per dataset for a given architecture and $50,220$ for the 9 dataset/model combinations.
5.5 Privacy Evaluation
5.5.1 Unlearning-Membership Inference Attack (U-MIA)
5.5.2 Unlearning-Likelihood Ratio Attack (U-LiRA)
The performance of a general MIA can be improved by considering, e.g., a per-sample attack such as LiRA [44, 45]. For any given point, we wish to determine whether the outputs from the unlearned models differ from those of models that have never seen the data point. To assess the attack robustly, we evaluate it across multiple models, using shadow models trained on various retain/forget sets. Specifically, we first train $n$ models based on $n$ splits of the training data. This train data is then split into 10 random retain and forget splits; hence, we unlearn a total of $10n$ models. We then perform hyper-parameter sweeps, similar to what we do in the original results and unlearn using the optimal hyper-parameters, except that we consider $\frac{n}{2}$ sweeps and conduct $200$ trials per sweep to determine the best hyper-parameters. In our setting, we set $n=64$.
TABLE III: Ranking by performance on Retention Deviation and Indiscernibility across datasets and architectures. We count the number of times each method appears in the Best Performers group (G1), Average performance group (G2) and Worst performers group (G3) (see §5). The final rank is computed based on the number of times the method appears in G1 with occurrences in G2 and G3 used to break ties if needed. If a method does not produce any usable models, it is assigned to a Failed group (F). Three methods appear in the top 3 for both performance measures: MSG (1st and 1st), CT (3rd and 1st) and KDE (3rd and 2nd).
Retention Deviation Indiscernibility Pareto Frontier
Rank Method G1 G2 G3 F Rank Method G1 G2 G3 F Rank Method Times on Frontier
1 FT 8 1 0 0 1 CT 9 0 0 0 1 CFW 4
1 MSG 8 1 0 0 1 MSG 9 0 0 0 2 CT 3
2 PRMQ 7 2 0 0 2 CFW 7 2 0 0 2 FT 3
3 CT 7 1 1 0 2 RNI 7 2 0 0 3 PRMQ 2
3 KDE 7 1 1 0 2 KDE 7 2 0 0 4 KDE 1
3 CFW 7 1 1 0 3 FT 6 3 0 0 4 FCS 1
4 FCS 6 3 0 0 3 PRMQ 6 3 0 0 4 SCRUB 1
4 SalUN 6 3 0 0 3 SalUN 6 3 0 0 4 SRL 1
5 NG+ 5 4 0 0 4 SRL 6 2 1 0 4 CF-k 1
5 SRL 5 4 0 0 5 NG+ 5 4 0 0 5 BT 0
6 SCRUB 4 3 1 1 5 FCS 5 4 0 0 5 GA 0
7 BT 2 7 0 0 6 SCRUB 5 3 0 1 5 MSG 0
7 RNI 2 7 0 0 7 BT 4 5 0 0 5 NG+ 0
8 CF-k 2 3 2 2 8 CF-k 1 2 4 2 5 RNI 0
9 IU 1 0 2 6 9 EU-k 1 2 2 4 5 SalUN 0
10 EU-k 0 5 0 4 10 GA 0 4 4 1 5 EU-k 0
11 GA 0 1 7 1 11 IU 0 0 3 6 5 IU 0
12 FF 0 0 0 9 12 FF 0 0 0 9 5 FF 0
5.6 Ranking the methods
@@
A challenge in comparing MU method performance comes from the potential proximity of the evaluation metrics. As a simple example, suppose we have four methods $U_{1},...,U_{4}$ with accuracies: 98%, 99%, 50%, 1%, respectively; if we simply rank the methods, the rank itself would not be representative of the fact that, e.g., $U_{1}$ and $U_{2}$ are much above $U_{3}$ and $U_{4}$. To enable distinctions based on proximities, we use Agglomerative Clustering and define cut-off points such that we obtain three clusters: (1) Best performers (G1), (2) Average performers (G2), and (3) Worst performers (G3). If a method does not produce ten usable models, one per original model, we assign it to the Failed group (F). For each method, we count the times it appears in the three groups (with nine being the maximum). To obtain a final ranking of the methods, we first rank them using the number of times they appear in the Best Performers group (G1); if ties occur, we break them with the Average Performers (G2) group. If ties persist, the Worst Performers (G3) group is the final tie-breaker. This method ensures a clear and fair ranking by considering each performance group in order of importance.
5.6.1 Pareto Frontier
We rank methods based on the number of times they lie on the Pareto Frontier. The Pareto frontier represents the set of methods that are non-dominated, that is, methods such that no other method performs better on both metrics simultaneously.
To determine this, we consider two metrics: (1) The Performance Retention Deviation metric to be minimized and (2) The Indiscernibility to be maximized.
To identify the Pareto frontier, we proceed as follows:
  1. 1.
    We gather the values of the two metrics for each method.
  2. 2.
    We sort the methods by the first metric in ascending order (since it needs to be minimized) and secondarily by the second metric in descending order (since it needs to be maximized).
  3. 3.
    We iterate through the sorted methods and retain only those not dominated by any other.
  4. 4.
    The remaining methods after this filtering step constituted the Pareto frontier, representing the optimal trade-off between the two metrics, where improving one metric would necessarily worsen the other.
5.7 Consistent Unforgettability
We question two aspects:
  1. 1.
    Are there images for which the U-MIA is consistently successful across model seeds?
  2. 2.
    Are there images that multiple methods fail to unlearn?
To answer these questions, we perform a 10-fold cross-validation U-MIA. Namely, given a model $f_{\theta}$ and two sets of equal size, one containing only the losses associated with images from the Forget set $\mathcal{D}_{f}$, and the other containing only losses associated with images from the Test set $\mathcal{D}_{T}$, we take the union of these sets and generate $10$ subset in direct sum. We then train $10$ Logistic Regression classifiers, each using $9$ subsets to predict the remaining. We keep track of each prediction made by the $10$ models, allowing us to get predictions for the entire Forget set. We apply each unlearning method to the Original model resulting from each of the $10$ initialization seeds. Since the forget set is the same across all experiments, we can now determine how many times each of the elements of $\mathcal{D}_{f}$’s membership is correctly inferred. From these, we can answer both questions, first by analyzing across seeds and then by analyzing across unlearning methods.
@@ 6 Main Results
We ranked the MU methods based on three criteria: Performance Retention Deviation, Indiscernibility, and the number of times they lie on the Pareto Frontier (Table III). We provide the performance table for CIFAR-100 using ResNet-18 (Table IV). Since the groups are obtained using a clustering algorithm, they do not convey information about the distance between groups; as such, we provide such information for FashionMNIST and CIFAR-100 (Table V). We gathered the Run-Time Efficiency of each method across datasets for the ResNet-18 architecture (Table VI) We provide we provide per dataset and architecture results in the Appendix.
TABLE IV: CIFAR-100 - ResNet18. CIFAR-100 provides the most visible comparison as there is a large gap in performance between the Retain Set and Test set, this leads to much larger RetDev scores. O and N represent the Original and Retrained models. The Test set MIA against the Original model disinguishes between the Test and Froget sets data points with $73\%$, accuracy while most unlearning methods reach MIA score close to random guessing ($50\%$).
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
Unlearner
BT 0.98 0.68 0.54 0.98 1.25 0.99 0.27 0.95 0.48 9.39
CF-k 1.00 0.83 0.56 1.00 1.53 1.02 0.55 0.73 0.63 5.91
CFW 0.98 0.43 0.43 0.98 0.79 0.78 0.44 1.00 0.50 6.17
CT 0.99 0.53 0.53 0.99 0.97 0.97 0.07 0.99 0.49 11.82
EU-k - - - - - - - - - -
FCS 0.98 0.54 0.55 0.98 0.99 1.01 0.04 0.92 0.54 3.02
FF - - - - - - - - - -
FT 0.98 0.55 0.54 0.98 1.02 0.98 0.05 0.99 0.50 5.16
GA 0.34 0.33 0.24 0.34 0.60 0.44 1.61 0.90 0.55 39.97
IU - - - - - - - - - -
KDE 0.99 0.52 0.51 0.99 0.95 0.94 0.11 0.99 0.50 3.98
MSG 0.91 0.38 0.38 0.91 0.69 0.69 0.71 1.00 0.50 4.49
NG+ 0.89 0.59 0.49 0.89 1.08 0.89 0.29 0.98 0.49 12.14
O 0.98 0.98 0.56 0.98 1.81 1.02 0.85 0.53 0.73 1.10
PRMQ 0.97 0.47 0.46 0.97 0.86 0.85 0.32 1.00 0.50 4.34
R 1.00 0.55 0.55 1.00 1.00 1.00 0.00 0.99 0.49 1.00
RNI 0.99 0.45 0.45 0.99 0.83 0.82 0.36 0.98 0.49 3.65
SCRUB 0.97 0.50 0.53 0.97 0.91 0.96 0.15 0.98 0.51 3.81
SRL 1.00 0.55 0.52 1.00 1.00 0.95 0.06 0.98 0.49 3.67
SalUN 0.98 0.49 0.51 0.98 0.91 0.93 0.18 0.99 0.49 10.66
TABLE V: Group Distances: Average scores per performance groups (G1: Best performers, G2: Average performers, G3: Worst performers). This allows for a finer-grained comparison of the groups. For instance, on CIFAR-100, with a ResNet18 architecture, and considering the RetDev, the average of the methods from the G3 group is $1.18$ away from the average RetDev of the G2 methods, which is $0.33$ away from G1. Namely, The methods from G3 are 3.58×3.58$\times$3.58 × further apart from the G2 group than G2 from G1
FashionMNIST Performance Retention Deviation (RetDev)
Architecture Worst G3 Mean G2 Mean G1 Mean Best
ResNet18 0.09 0.08 0.05 0.02 0.01
TinyViT 0.17 0.17 0.08 0.04 0.02
Indiscernibility
Architecture Worst G3 Mean G2 Mean G1 Mean Best
ResNet18 0.87 0.87 0.92 0.98 0.99
TinyViT 0.93 0.93 0.98 0.99 1.00
CIFAR-100 Performance Retention Deviation (RetDev)
Architecture Worst G3 Mean G2 Mean G1 Mean Best
ResNet18 1.61 1.61 0.42 0.09 0.04
TinyViT 2.15 2.15 0.44 0.13 0.10
Indiscernibility
Architecture Worst G3 Mean G2 Mean G1 Mean Best
ResNet18 0.73 0.73 0.92 0.99 1.00
TinyViT 0.87 0.88 0.94 0.99 1.00
6.1 On the reliability of baselines.
@@
The commonly used baseline FT trains the original model only on the retain set for several epochs to enable the model to forget information about the forget set. In our evaluation, FT performs best based on the Retention Deviation and is ranked third based on Indiscernibility (Table III). The latter observation may come from the fact that FT does not explicitly unlearn the forget set or perturb the model parameters. Based on these results, we conclude that it is a reasonable baseline against which to evaluate (Tables XVI, XVII). We, however, remark that since the mechanism underlying FT (training on the Retain set to maintain performance) is common to many other methods, these methods may inherit its susceptibility to MIA. Another common baseline, GA, which performs gradient ascent on the forget set, performs poorly across both metrics. Its more recent variation, NG+, which uses an additional retain set correction, ranks fifth for both metrics, making it a more suitable baseline.
@@
TABLE VI: Run Time Efficiency on ResNet for the top performing methods. CT is the fastest on average, and MSG runs up to 5x faster than naive retraining.
@@
6.2 On the reliability of newly proposed unlearning methods.
@@
MSG obtained the first rank in both performance retention deviation and indiscernibility (Table III). Its unique approach identifies the parameters in the Convolutional layers that most contribute to the information being forgotten. This strategy, differing from FT, allows MSG to retain performance from the Retain set while modifying the weights more relevant to the Forget set. PRMQ ranks second in performance retention, making it one of the top performers. However, it suffers from the same lower performance in Indiscernibility as FT. PRMQ, however, does not leverage the Forget set; instead, it performs a form of knowledge distillation by attempting to reproduce the results of the original models on the Retain set. During the pruning phase, the weights for the MLP and Convolutional layers are reinitialized. CT ranks third in Performance Retention and first in Indiscernibility. It is interesting to note that CT and MSG are consistently among the top performers in Indiscernibility, whereas FT performs poorly (Table III).
Figure 2: Consistent Unforgettability on CIFAR-100: For each image in the Forget set, we determine the number of models for which the weak MIA is successful for this image. Even for the Retrained model R, $10\%$ of the Forget Set is detected as such across all 10 random model initializations for the ResNet18 model. Nonetheless, for SCRUB close to $17.5\%$, forget sets are never detected as part of the Forget set through the U-MIA. Furthermore, the patterns across architectures differ greatly, especially for SCRUB.
6.3 On the robustness across architectures.
@@
A critical aspect of a good MU method is its ability to generalize across various DNN architectures. We conducted experiments with both ResNet18 and TinyViT (Tables XVI, Table XVII). Despite being tailored to Convolutional Neural Networks (CNN) models, methods such as CT still perform competitively when applied to Vision Transformers (Table III). While proposed for CNN layers, methods such as CT and MSG work well on Vision Transformer as one can leverage the 2D Convolutions used in Positional Encoders. We provide additional details on the ranking based on architectures in Appendix B.
6.4 On the speed of the unlearning methods.
@@
MSG, the best MU candidate, runs  7.6x faster than retraining from scratch on UTKFace and  3.3x on FashionMNIST (Table VI). On average, MSG runs  5.3x faster than retraining. The fastest method is CT, which achieves a speedup of  17.5x compared to Retraining on CIFAR-10 and is, on average, 10.59×10.59$\times$10.59 × faster than retraining across all datasets. This speedup stems from its simple approach: transposing the convolutional layer weights.
6.5 On the Consistent Unforgettability.
We evaluate the concept of Consistent Unforgettability on CIFAR-100 (Figure 2). For each method, we count the number of times the membership of a given point of $\mathcal{D}_{F}$ is correctly inferred when compared to points of the $\mathcal{D}_{T}$ (Figure 2). First, we note that even for the Retrained model, which by design has never seen data points from the Forget set $\mathcal{D}_{f}$, $10\%$ of the Forget set images are identified as belonging to the Forget set across all $10$ model seeds. Furthermore, $59.6\%$ and $65.2\%$ of the Forget Set images are correctly predicted to belong to the Forget Set for at least five seeds. Furthermore, Figure 2 shows despite using the same hyper-parameters and the same data splits, the performance of an unlearning method can vary across seeds and architecture. This calls for comparing unlearning models across original model initialization seeds to evaluate the reliability of such a method. Additionally, we observed data points in the forgot set of CIFAR-100 that are correctly identified across all ten initialization seeds by 13 methods out of 18; the 13 methods might differ from one image to the other. Nonetheless, this shows shared sets of difficult images to forget.